OpenStudy (anonymous):

Two is thirty-one, divided by the sum of a number and four. Find the number. A. 11.5 B. –15.5 C. 13.5 D. none of the above

OpenStudy (anonymous):

\[2 = 31/(x+4)\]

OpenStudy (anonymous):

2x+8=31 2x=23 x=23/2

OpenStudy (anonymous):

so, A

OpenStudy (texaschic101):

your not explaining enough about how you are getting your answers